# Article Category

### v2

@@ 6.3 EVALUATION OF SAFEGUARDS
The impeded learning performance was visible in the quadrotor environment as a drop in policy quality and convergence speed for boundary projection, as the agents were learning very slowly or completely stalled for several individual runs. We made the same observation for the ray mask to a lesser extent. In contrast, the convergent runs in the pendulum environment showed barely any degradation compared to unsafe training due to the simplicity of the environment. Furthermore, ray masking improved performance in the energy system environment substantially, which can be attributed to the guided exploration.
@@
For the boundary projection, the regularisation term alleviated most issues as performance was on par with unsafe training. The observed reduction in non-convergent runs suggests that regularisation improves convergence. However, the fact that non-convergent runs persisted rather than being eliminated indicates that the regularisation coefficient may be too small. The observation that non-convergent runs involve more safeguarding interventions than convergent ones supports this assumption. The impact of regularisation is less pronounced for ray masking, as it improved the convergence speed in the quadrotor environment, but not to the level of unsafe training. We attribute the negligible effect to the fact that regularising the ray mask constantly introduces a gradient component towards the centre. In contrast, regularisation only influences policy updates when actions are unsafe for boundary projection.
@@
Better performance from a passthrough gradient is possible since a robust control invariant state set captures most of the optimal actions in balancing tasks, which retains the gradient correctness while eliminating the gradient decrease in the mapping direction. Since the unaltered ray mask is almost optimal in the pendulum environment, performance increases are only visible in the quadrotor environment, where the gap to unsafe training is closed. The non-convergence of some passthrough runs could be due to the effectively increased learning rate in the balancing tasks. In the energy system task, not all reward-maximising actions are safe, such that the gradients of the passthrough safeguard can be wrong and therefore stall learning.
@@
The hyperbolic ray mask produces a similar mapping distance to boundary projection due to the hyperbolic tangent function, leading us to expect comparable performance. Unlike boundary projection, the hyperbolic map ensures that a gradient is always available. However, for unsafe actions, this gradient remains small. We observe marginally more stable convergence but significantly lower policy quality than boundary projection in the balancing tasks. This result is unexpected and may be attributed to the diminished gradient in the mapping direction, as indicated by the frequent safeguarding interventions. The performance of the regularised, hyperbolic ray mask supports this statement, as it achieved the best performance in the quadrotor environment and converged in all ten runs. The large confidence interval and poor mean performance in the pendulum environment were attributed to a single outlier, which was convergent but significantly slower than all other runs.
6.4 COMPARISON OF SAFE CENTRE APPROXIMATIONS
@@
We also compare safe centre approximations for the ray mask. Since the orthogonal approximation only applied to unsafe actions, safe actions were not mapped. In the pendulum task, the one-dimensional action space allows for exact safe centre approximations, which condenses the comparison to rarer interventions by the orthogonal approximation versus the smoother map of the zonotopic approximation. The continuous map provided by the zonotopic approximation produced superior final policies and converged faster. The low number of steps of the orthogonal approximation in the quadrotor task was an artefact of the worse policy, as seen in the full figures in Section˜7.5. In the energy system, the orthogonal approach converges faster initially but to a worse policy. Therefore, smooth safeguards appear more critical than rare interventions for learning.
@@
Pendulum Zonotopic 27392 [21241, 32768] -8 [-8, -8] 0 / 10
Orthogonal 30208 [18432, 39424] -8 [-8, -8] 0 / 10
Quadrotor Zonotopic 67148 [30808, 101287] -251 [-306, -194] 0 / 10
Orthogonal 31372 [5476, 57241] -432 [-482, -371] 0 / 10
Energy System Zonotopic 709280 [579898, 829873] -8793 [-12852, -3797] 0 / 10
@@
6.5 COMPARISON OF WALL CLOCK TIME
To estimate computational overhead, we compare the relative wall-clock time of safeguarded training with its unsafe counterpart in Table˜6. We measured the wall clock time for 10000 steps in the pendulum environment. We observed at least a four-fold increase in computation time with boundary projection. Ray masking took almost double the time of boundary projection, which we trace to the increased computational complexity of its optimisation problems due to the induced safe action set. SHAC produced around a quarter of the additional computational overhead, as it must maintain the computational graph for backpropagation. The wall clock time poses a significant downside, although a custom, more efficient implementation could mitigate the effects. Moreover, safeguarding via a ray mask is significantly cheaper for explicit safe action sets since the safe centre is provided and does not require costly approximation. However, pre-computing the safe state or action set may not be possible depending on the task, which could further increase the computation needed per training iteration.
Table 6: Relative wall clock time of the different safeguards for 10000 steps in the pendulum environment compared to their unsafe versions.
@@ 7 LIMITATIONS AND CONCLUSION
This work demonstrates the fundamental applicability and effectiveness of safeguards for analytic gradient-based reinforcement learning, unlocking its usage for safely training agents in simulations before deploying them in safety-critical applications. While we showcased the possibility of achieving performance on par or exceeding unsafe training, success depends on the quality and representation of the safe set.
While we utilised zonotopes, the safeguards presented are not limited to this set representation. The only limitation of the representation is the star-shapedness for the ray mask and the ability to solve the relevant containment problems in an efficient and differentiable manner. The general trade-off in the choice of set representation is achieving a tight approximation of the true safe set versus cheap containment problems. Tighter approximations allow for larger time steps due to the larger safe sets. In contrast, simple representations may allow a closed-form solution and compensate by decreasing the time step size.
A limitation of the zonotope representation is its symmetry. Symmetric safe set representations can unnecessarily limit the size of the safe set. The size limitations are evident in an obstacle avoidance scenario, where the closest distance to an obstacle limits the safe set size direction towards and away from the obstacle.
Moreover, solving the containment problems with CVXPY allows for rapid prototyping but may not offer optimal performance compared to custom solvers and formulations, which could decrease the substantial overhead. The group behind CVXPY recently addressed this issue with a parallel interior point solver [73] and CVXPYgen, which generates a custom solver in C [74], which should achieve substantial speed-ups in future work.
In general, safeguarding for the sole sake of efficiency requires either an informative, safe set or an expensive simulation since the sample efficiency gains strongly depend on the quality of the safe set. In contrast, the computational overhead depends only on the representation and dimensionality.
@@
7.1 ENVIRONMENT DESCRIPTIONS
All environments share some characteristics. The feasible state and action sets are axis-aligned boxes. The feasible action set is also of unit length. The transition distribution is given as a first-order ordinary differential equation with additive, bounded noise, which we integrate using an Euler scheme
$P_{f}(s_{i+1}|s_{i},a_{i})=s_{i}+dt\dot{s}_{i}+\mathcal{W}$ (30)
with the noise zonotope $\mathcal{W}$. For the energy system, we utilise an explicit Euler scheme, whereas for the pendulum and quadrotor, we use a semi-implicit Euler scheme
$P_{f}(s_{i+1}|s_{i},a_{i})=\begin{bmatrix}s_{i}\\ \dot{s}_{i}\end{bmatrix}+dt\begin{bmatrix}\dot{s}_{i+1}\\ \ddot{s}_{i}\end{bmatrix}+\mathcal{W}\,,$ (31)
where we exploit the form of our states.
Pendulum The environment possesses a feasible state set $S=\left[-\pi,\pi\right]\times\left[-8,8\right]$ with the state $s=\begin{bmatrix}\theta&\dot{\theta}\end{bmatrix}^{T}$ representing the angle $\theta$ and the angular velocity $\dot{\theta}$. The feasible action set is $\mathcal{A}=[-1,1]$ with the action $a$ representing the torque. The time derivative of the state is
@@ 7.1 ENVIRONMENT DESCRIPTIONS
$r(s,a)=-\theta^{2}-\frac{\dot{\theta}^{2}}{10}-\frac{a^{2}}{100}\,.$ (33)
The reward function encodes the goal of balancing the pendulum upright. We define the safety constraints as the part of the state space from which the controller can maintain balance, effectively limiting the velocity and angle close to the upright position. We induce a safe action set from a robust control invariant (RCI) state set, which we obtain by the method in Schäfer et al. [46]. Since the RCI set accounts for errors in the Taylor expansion, we linearise the dynamics but retain the original noise set $\mathcal{W}$ without expansion.
Quadrotor The environment possesses a feasible state set
@@
×$\times[-1.0,1.0]\times\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$
@@
with the torque magnitude $c_{0}$, the roll angle magnitude $c_{1}$, and the PID gains $pd_{0-2}$. The noise zonotope is $\mathcal{W}=\left<\textbf{0},I\begin{bmatrix}0.1&0.1&0&0&0&0\end{bmatrix}^{T}\right>$, where we denote the diagonalisation of the vector by the multiplication with the identity. The reward is
$\begin{split}r(s,a)&=-2.5\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}\\ &-\frac{r+\dot{x}+\dot{y}+\dot{r}}{10}-\frac{(a_{0}c_{0}+g)^{2}}{50}\\ &-\frac{(a_{1}c_{1})^{2}}{100}\,,\end{split}$ (36)
where $x_{0},y_{0}$ encode the position of the quadrotor at reset. The reward function encodes balancing the quadrotor around its spawning location. Similar to the pendulum, the safety constraints prevent escalation of the position and velocity, such that balancing always remains possible within the finite time horizon of the problem. Again, we induce a safe action set from an RCI set using the same approach as the pendulum.
Energy Management System The system has the feasible state set $\mathcal{S}=[0,10]\times[18,24]\times[10,100]$ with the state $s=\begin{bmatrix}e&\vartheta^{\text{in}}&\vartheta^{\text{ret}}\end{bmatrix}$, where $e$ is the charge of the battery, $\vartheta_{\text{in}}$ is the indoor temperature of the building, and $\vartheta_{\text{ret}}$ is the return temperature of the floor heating system. The feasible action set is $\mathcal{A}=[-1,1]^{2}$, representing the power set point for the battery $a_{0}$ and the heat pump $a_{1}$. The time derivative of the state is
@@
where the computation of the coefficients $c_{1-3}$ is detailed in [75].We assume a constant coefficient of performance but model the outdoor temperature and the resulting coefficient as a random variable bounded by the noise zonotope
@@
where $p^{\ell}$ is the inflexible load of the building, $p^{\text{PV}}$ is the output of the photovoltaic generator, $\phi$ is the electricity price, and $\vartheta^{\text{set}}$ is the desired indoor temperature. The reward function encodes the goal of minimising energy expenditure while maintaining room temperature. To facilitate the task the observation $o_{t}=\begin{bmatrix}s_{t}&\vartheta^{\text{out}}_{[t:t+H]}&p^{\text{PV}}_{[t:t+H]}&p^{\ell}_{[t:t+H]}&\phi_{[t:t+H]}\end{bmatrix}$ with the outdoor temperature $\vartheta^{\text{out}}$ additionally contains forecasts besides current measurements. We choose $H=$, resulting in 272727 observations. The safe state set is the subset of the feasible state set from which the agent can recover using the full action range within one time step. Since the state transitions are linear, we do not require an approximation.
7.2 LEARNING CURVES OF ALL LEARNING ALGORITHMS IN NON-SAFEGUARDED TRAINING
7.3 LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH BOUNDARY PROJECTION
7.4 LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH RAY MASK
Figure 9: Learning curves SHAC, SAC, and PPO with ray mask.
7.5 LEARNING CURVES OF BOTH APPROXIMATIONS
7.6 PROOF OF LEMMA 1: JACOBIAN OF BOUNDARY PROJECTION
@@
Equation˜12 becomes
@@
with safe action set $\mathcal{A}_{s}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right>$. Translating Equation˜40 to canonical, quadratic form yields the optimisation variable $z=\begin{bmatrix}a_{s}\\ \gamma\end{bmatrix}\in\mathbb{R}^{d+n}$ and the problem
@@
$K$ $=\begin{bmatrix}\textbf{0}&I_{n}\end{bmatrix}\in\mathbb{R}^{n\times(d+n)}$ (46)
$h$ $=\textbf{1}\in\mathbb{R}^{n}\,,$ (47)
where the subscript of the identity denotes its size, and bold scalars are matrices of appropriate size filled with the scalar. To compute the rank of the Jacobian, we start from the differentials of the KKT conditions [76, Eq. 6]
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&Kz_{D}^{*}-h_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\mathop{}\!\mathrm{d}z\\ \mathop{}\!\mathrm{d}\lambda\\ \mathop{}\!\mathrm{d}\nu\end{bmatrix}=$ (48)
where the subscript $D$ denotes a diagonalised vector, the superscript * optimal values, $\nu\in\mathbb{R}^{d}$ the dual variables on the equality constraints, $\lambda\in\mathbb{R}^{n}$ the dual variables on the inequality constraints, and $\mathop{}\!\mathrm{d}$ a differential. To obtain the Jacobian with respect to the action, we substitute $\mathop{}\!\mathrm{d}q=\frac{\mathop{}\!\mathrm{d}q}{\mathop{}\!\mathrm{d}a}\mathop{}\!\mathrm{d}a=-2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\mathop{}\!\mathrm{d}a$ and all other differential terms with zero, leaving
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&Kz_{D}^{*}-h_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\mathop{}\!\mathrm{d}z}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}\end{bmatrix}=\begin{bmatrix}2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\\ \textbf{0}\\ \textbf{0}\end{bmatrix}\,.$ (49)
We insert the variables and describe the system in expanded form as
$\begin{bmatrix}2I_{d}&\textbf{0}&\textbf{0}&I_{d}\\ \textbf{0}&\textbf{0}&I_{n}&-G_{\mathcal{A}_{s}}^{T}\\ \textbf{0}&\lambda_{D}^{*}&\gamma_{D}^{*}-I_{n}&\textbf{0}\\ I_{d}&-G_{\mathcal{A}_{s}}&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}\\ \end{bmatrix}=\begin{bmatrix}2I_{d}\\ \textbf{0}\\ \textbf{0}\\ \textbf{0}\\ \end{bmatrix}\,.$ (50)

### v3

@@ 6.C EVALUATION OF SAFEGUARDS
Boundary projection with regularisation alleviated most issues of the unaltered variant, as performance was on par with unsafe training. The observed reduction in non-convergent runs suggests that regularisation improves convergence. However, the fact that non-convergent runs persisted rather than being eliminated, indicates that the regularisation coefficient may be too small. The observation that non-convergent runs involve more safeguarding interventions than convergent ones supports this assumption. Due to time constraints, we could not run additional experiments with an increased regularisation coefficient.
Ray masking with regularisation had less changes, as it improved the convergence speed in the quadrotor environment, but not to the level of unsafe training. We attribute the negligible effect to the fact that regularising the ray mask constantly introduces a gradient towards the centre. In contrast, regularisation only influences policy updates when actions are unsafe for boundary projection.
@@
Ray masking with a passthrough gradient can improve performance since a robust control invariant state set captures most of the optimal actions in balancing tasks, which retains the gradient correctness while eliminating the gradient decrease in the mapping direction. Since the unaltered ray mask is almost optimal in the pendulum environment, performance increases are only visible in the quadrotor environment, where the gap to unsafe training is closed. The non-convergence of some runs in the balancing tasks could be due to the effectively increased learning rate, as the gradient is no longer diminished by the safeguard. In the energy system task, not all reward-maximising actions are safe, such that the gradients of this safeguard can be wrong and therefore stall learning.
@@
The hyperbolic ray mask produces a similar mapping distance to boundary projection due to the hyperbolic tangent function, leading us to expect comparable performance. Unlike boundary projection, the hyperbolic map ensures that a gradient is always available. However, for unsafe actions, this gradient remains small. We observe marginally more stable convergence but significantly lower policy quality than boundary projection in the balancing tasks. This result is unexpected and may be attributed to the diminished gradient in the mapping direction, as indicated by the frequent safeguarding interventions. The performance of the regularised, hyperbolic ray mask supports this statement, as it achieved the best performance in the quadrotor environment and converged in all ten runs. The large confidence interval and poor mean performance in the pendulum environment were attributed to a single outlier, which converged significantly slower than all other runs.
6.D COMPARISON OF SAFE CENTRE APPROXIMATIONS
@@
We also compare safe centre approximations for the ray mask, where we found that the zonotopic approximation results in superior final policies and faster convergence, see Table˜5. Since the orthogonal approximation only applied to unsafe actions, safe actions were not mapped. In the pendulum task, the one-dimensional action space allows for exact safe centre approximations, which condenses the comparison to rarer interventions by the orthogonal approximation versus the smoother map of the zonotopic approximation. The continuous map provided by the zonotopic approximation produced superior final policies and converged faster. The low number of steps of the orthogonal approximation in the quadrotor task was an artefact of the worse policy, as seen in the learning curves in Appendix˜G. In the energy system, the orthogonal approach converged faster initially, but to a worse policy. Therefore, smooth safeguards appear more critical than rare interventions for learning.
@@
Pendulum Zonotopic 27392 [021241, 032768] -8 [0000-8, 0000-8] 0 / 10
Orthogonal 30208 [018432, 039424] -8 [0000-8, 0000-8] 0 / 10
Quadrotor Zonotopic 67148 [030808, 101287] -251 [00-306, 00-194] 0 / 10
Orthogonal 31372 [005476, 057241] -432 [00-482, 00-371] 0 / 10
Energy System Zonotopic 709280 [579898, 829873] -8793 [-12852, 0-3797] 0 / 10
@@
6.E COMPARISON OF COMPUTATION TIME
The relative computation time of safeguarded training is compared to its unsafe counterpart in Table˜6 to estimate computational overhead. For this purpose, the computation time was measured over 10,000 steps in the pendulum environment. The results show at least a four-fold increase in computation time when boundary projection is applied. Ray masking took almost double the time of boundary projection, which we trace to the increased computational complexity of its optimisation problems, due to the derived safe action set. SHAC produced around a quarter of the additional computational overhead, as it must maintain the computational graph for backpropagation. The increased computation time poses a significant downside, although a custom, more efficient implementation could mitigate the effects. Moreover, safeguarding via a ray mask is significantly cheaper for specified safe action sets, since the safe centre is provided. However, pre-computing the safe state or action set may not be possible depending on the task, which could further increase the computation needed per training iteration.
Table 6: Relative computation time of the different safeguards for 10,000 steps in the pendulum environment compared to their unsafe versions.
@@ 7 LIMITATIONS AND CONCLUSION
This work demonstrates the fundamental applicability and effectiveness of safeguards for analytic gradient-based reinforcement learning, unlocking its usage for safely training agents in simulations before deploying them in safety-critical applications. While we showcased the possibility of achieving performance on par with or exceeding unsafe training, success depends on the quality and representation of the safe set.
While we utilised zonotopes, the safeguards presented are not limited to this set representation. The only limitation of the representation is the star-shapedness of the ray mask and the ability to solve the relevant containment problems in an efficient and differentiable manner. The general trade-off in the choice of set representation is achieving a tight approximation of the true safe set versus computationally cheap containment problems. A limitation of the chosen zonotope representation is its inherent symmetry, which can unnecessarily restrict the safe set. This restriction becomes apparent near the boundary of $\mathcal{A}$, where the minimum distance to the safe set boundary constrains the extent of the set both towards and away from it.
Moreover, using CVXPY allows for rapid prototyping but may not offer optimal performance compared to custom solvers and formulations, which could decrease the substantial overhead of safeguarding. The group behind CVXPY recently addressed this issue by a parallel interior point solver [77] and CVXPYgen [78], which generates a custom solver in C.
In general, safeguarding for the sole sake of efficiency requires either an informative, safe set or an expensive simulation since the sample efficiency gains strongly depend on the quality of the safe set. In contrast, the computational overhead depends only on the representation and dimensionality of the safe action set.
@@
A ENVIRONMENT DESCRIPTIONS
All environments share some characteristics: the feasible state and action sets are axis-aligned boxes; the feasible action set is of unit length; and the dynamics is given as a first-order ordinary differential equation, which we integrate using an Euler scheme. The non-determinism of the system is encapsulated by additive bounded noise. This yields the transition function
$s_{i+1}=s_{i}+dt\left(\dot{s}_{i}+w_{i}\right)$ (33)
with the discrete time step size $dt$ and the noise sample $w_{i}\in\mathcal{W}$, where $\mathcal{W}$ is a zonotope. For the energy system, we utilise an explicit Euler scheme, whereas for the pendulum and quadrotor, we use a semi-implicit Euler scheme
$s_{i+1}=\begin{bmatrix}p_{i}\\ \dot{p}_{i}\end{bmatrix}+dt\left(\begin{bmatrix}\dot{p}_{i+1}\\ \ddot{p}_{i}\end{bmatrix}+w_{i}\right)\,$, (34)
where we exploit the form of our state $s=\begin{bmatrix}p\\ \dot{p}\end{bmatrix}$, which consists only of coordinates $p$ and their respective velocities $\dot{p}$. We choose a semi-implicit Euler integrator, since it is symplectic.
Pendulum This environment possesses a feasible state set $S=\left[-\pi,\pi\right]\times\left[-8,8\right]$ with the state $s=\begin{bmatrix}\theta&\dot{\theta}\end{bmatrix}^{T}$ representing the angle $\theta$ and the angular velocity $\dot{\theta}$. The feasible action set has one dimension with the action $a$ representing the torque. The dynamics is
@@ A ENVIRONMENT DESCRIPTIONS
$r(s,a)=-\theta^{2}-\frac{\dot{\theta}^{2}}{10}-\frac{a^{2}}{100}$ (36)
and encodes the goal of balancing the pendulum upright. Colloquially, we define safety as the part of the state space from which the controller can maintain balance without the pendulum falling. Formally, we derive a safe action set from a robust control invariant (RCI) state set, which we obtain by the method in Schäfer et al. [46].
Quadrotor This environment possesses a feasible state set
@@
$\times[-1.0,1.0]\times\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$
with the state $s=\begin{bmatrix}x&y&r&\dot{x}&\dot{y}&\dot{r}\end{bmatrix}^{T}$ representing the quadrotor position $(x,y)$, roll $r$, and their respective velocities $(\dot{x},\dot{y},\dot{r})$. The feasible action set has two dimensions with the thrust $a_{0}$ and roll angle $a_{1}$. The dynamics is
with the torque magnitude $c_{0}$, the roll angle magnitude $c_{1}$, and the PID gains $pd_{0-2}$. The noise zonotope is $\mathcal{W}=\left<\textbf{0},\begin{bmatrix}0.1&0.1&0&0&0&0\end{bmatrix}_{D}\right$>. The reward is
@@
where $x_{0},y_{0}$ encode the initial position of the quadrotor. The reward function encodes balancing the quadrotor around its initial location. Again, we derive a safe action set from an RCI set to obtain the set of safe actions.
Energy Management System This system has the feasible state set $\mathcal{S}=[0,10]\times[18,24]\times[10,100]$ with the state $s=\begin{bmatrix}e&\vartheta^{\text{in}}&\vartheta^{\text{ret}}\end{bmatrix}^{T}$, where $e$ is the charge of the battery, $\vartheta_{\text{in}}$ is the indoor temperature of the building, and $\vartheta_{\text{ret}}$ is the return temperature of the floor heating system. The feasible action set has two dimensions, representing the power set point for the battery $a_{0}$ and the heat pump $a_{1}$. The dynamics is
@@
where the computation of the coefficients $c_{1-3}$ is detailed in [79, Eq. 2.17]. The noise zonotope is
@@
where $p^{\ell}$ is the inflexible load of the building, $p^{\text{PV}}$ is the output of the photovoltaic generator, $\phi$ is the electricity price, and $\vartheta^{\text{set}}$ is the desired indoor temperature. The reward function encodes the goal of minimising energy consumption while maintaining room temperature. To facilitate the task, the observation $o_{t}=\begin{bmatrix}s_{t}&\vartheta^{\text{out}}_{[t:t+H]}&p^{\text{PV}}_{[t:t+H]}&p^{\ell}_{[t:t+H]}&\phi_{[t:t+H]}\end{bmatrix}^{T}$ includes the outdoor temperature $\vartheta^{\text{out}}$, current measurements, and forecasts, where we use the slicing notation $x_{\left[i:j\right]}=\begin{bmatrix}x_{i}&\dots&x_{j-1}\end{bmatrix}^{T}$. We choose $H=5$, resulting in $23$ observations. The safe state set is the feasible state set.
B PROOF OF THEOREM 1: JACOBIAN OF BOUNDARY PROJECTION
@@
We investigate the rank of the Jacobian of boundary projection Equation˜12, namely $\text{rank}(\frac{\partial g_{\text{BP}}(a)}{\partial a})=\text{rank}(\frac{\partial a_{s}}{\partial a})$, by utilising the differentials of the KKT conditions of a canonical, quadratic program [80, Eq. 6]
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&(Kz^{*}-h)_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\mathop{}\!\mathrm{d}z\\ \mathop{}\!\mathrm{d}\lambda\\ \mathop{}\!\mathrm{d}\nu\end{bmatrix}$= (43)
where the superscript * denotes optimal values, bold scalars a constant matrix of suitable size with all entries equal to the scalar, $\nu\in\mathbb{R}^{d}$ the dual variables on the equality constraints, $\lambda\in\mathbb{R}^{2n}$ the dual variables on the inequality constraints, and $\mathop{}\!\mathrm{d}$ a differential. Under the assumptions in Theorem˜1, Equation˜12 is
@@
which we reformulate in canonical, quadratic form
@@
$K$ $=\begin{bmatrix}\textbf{0}&I_{n}\\ \textbf{0}&-I_{n}\end{bmatrix}\in\mathbb{R}^{2n\times(d+n)}$ (51)
$h$ $=\textbf{1}\in\mathbb{R}^{2n}\,$, (52)
[…]
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&(Kz^{*}-h)_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\partial z}{\partial a}\\ \frac{\partial\lambda}{\partial a}\\ \frac{\partial\nu}{\partial a}\end{bmatrix}=\begin{bmatrix}2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\\ \textbf{0}\\ \textbf{0}\end{bmatrix}\,$. (53)
We insert Equations˜46, 47, 49, 51 and 52, such that $z^{*}=\begin{bmatrix}a_{s}^{*}&\gamma^{*}\end{bmatrix}^{T}$ and $\frac{\partial z}{\partial a}=\begin{bmatrix}\frac{\partial a_{s}}{\partial a}&\frac{\partial\gamma}{\partial a}\end{bmatrix}^{T}$, which expands the system into
$\begin{bmatrix}2I_{d}&\textbf{0}&\textbf{0}&I_{d}\\ \textbf{0}&\textbf{0}&{\begin{bmatrix}I_{n}&-I_{n}\end{bmatrix}}&-G_{\mathcal{A}_{s}}^{T}\\ \textbf{0}&\lambda_{D}^{*}\begin{bmatrix}I_{n}\\ -I_{n}\end{bmatrix}&\begin{bmatrix}\gamma_{D}^{*}-I_{n}&\textbf{0}\\ \textbf{0}&-\gamma_{D}^{*}-I_{n}\end{bmatrix}&\textbf{0}\\ I_{d}&-G_{\mathcal{A}_{s}}&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\partial a_{s}}{\partial a}\\ \frac{\partial\gamma}{\partial a}\\ \frac{\partial\lambda}{\partial a}\\ \frac{\partial\nu}{\partial a}\\ \end{bmatrix}$
$\qquad=\begin{bmatrix}2I_{d}\\ \textbf{0}\\ \textbf{0}\\ \textbf{0}\\ \end{bmatrix}\,$. (54)